A bag contains 6 red marbles, 9 blue marbles, and 5 green marbles. You withdrew one marble,replace it, and then withdraw another marble. what is the probability that you do NOT pick two green marbles?
Answer by venugopalramana(3286)

A bag contains 6 red marbles, 9 blue marbles, and 5 green marbles. You withdrew one marble,replace it, and then withdraw another marble. what is the probability that you do NOT pick two green marbles?
in such problems it is easier to find probability of an event happening and subtract it from total probability of 1 to get probability of event not happening.
SO LET US CONSIDER PROBABILITY OF DRAWING 2 GREEN MARBLES.
THERE ARE 6 RED+9 BLUE+5 GREEN MARBLES=20 MARBLES.
PROBABILITY OF GETTING A GREEN MARBLE ON FIRST ATTEMPT=5/20=1/4
PROBABILITY OF GETTING A GREEN MARBLE ON SECOND ATTEMPT=
=PROBABILITY OF GETTING A GREEN MARBLE ON FIRST ATTEMPT AS THE MARBLE IS REPLACED =5/20=1/4
HENCE PROBABILITY OF DRAWING 2 GREEN MARBLES = (1/4)8(1/4)=1/16
HENCE PROBABILITY OF NOT DRAWING 2 GREEN MARBLES =1-1/16=15/16